### v1

# From Pixels to Policies: Reinforcing Spatial Reasoning in Language Models for Content-Aware Layout Design

Abstract
We introduce LaySPA, a reinforcement learning framework that equips large language models (LLMs) with explicit and interpretable spatial reasoning for content-aware graphic layout design. LaySPA addresses two key challenges: LLMs’ limited spatial reasoning and the lack of opacity in design decision making. Instead of operating at the pixel level, we reformulate layout design as a policy learning problem over a structured textual spatial environment that explicitly encodes canvas geometry, element attributes, and inter-element relationships. LaySPA produces dual-level outputs comprising interpretable reasoning traces and structured layout specifications, enabling transparent and controllable design decision making. Layout design policy is optimized via a multi-objective spatial critique that decomposes layout quality into geometric validity, relational coherence, and aesthetic consistency, and is trained using relative group optimization to stabilize learning in open-ended design spaces. Experiments demonstrate that LaySPA improves structural validity and visual quality, outperforming larger proprietary LLMs and achieving performance comparable to specialized SOTA layout generators while requiring fewer annotated samples and reduced latency.
[…]
From Pixels to Policies: Reinforcing Spatial Reasoning in Language Models for Content-Aware Layout Design
Sha Li1, Stefano Petrangeli2, Yu Shen2, Xiang Chen2 1Virginia Tech 2Adobe Research
1 Introduction
Large Language Models (LLMs) have demonstrated strong capabilities in structured reasoning (Zhang et al., 2025), planning (Ferrag et al., 2025) and decision making (Zhai et al., 2025). These advances extend beyond natural language understanding and generation to domains requiring multi-step reasoning (Plaat et al., 2025), interaction (Chen et al., 2025), and symbolic manipulation, positioning LLMs as a promising substrate for design-oriented tasks that demand explicit relational and structural reasoning (Wu et al., 2025b). Within this context, automatic content-aware graphic layout design (Zheng et al., 2019) emerges as a compelling testbed for LLM-centric spatial reasoning, with applications spanning advertising (Hsu et al., 2023), web interfaces (Wang et al., 2025), and poster design (Chai et al., 2023a), all of which require coordinated placement of multiple elements under intertwined semantic, geometric, and aesthetic constraints.
((a))
((b))
((c))
Figure 1: Comparison between a human-designed poster and a GPT-5–generated layout, illustrating common spatial reasoning failures in LLM-based layout generation.
Prior approaches typically formulate layout design as coordinate regression or image-conditioned synthesis, using diffusion models (Chai et al., 2023b; Zhang et al., 2023) or autoregressive generators (Horita et al., 2024). While effective at producing visually plausible layouts, these methods rely heavily on large annotated datasets and offer limited interpretability: spatial decisions are entangled with perception in dense latent representations, often prioritizing appearance over structural principles. In contrast, layout design requires explicit global spatial coherence, multi-object relational reasoning and semantic–functional coupling (e.g., hierarchy, alignment, decoration), and generalizability to novel element configurations or constraints, which distinguishes it from pixel-level image generation and expose the limitations of purely perceptual modeling.
LLMs provide a complementary perspective as they encode prior knowledge of design principles (Feng et al., 2023), support symbolic reasoning (Wu et al., 2023) and produce interpretable outputs. Early attempts apply LLMs to layout generation via structured serializations using in-context learning (Horita et al., 2024), chain-of-thought reasoning (Wei et al., 2022), or prompt engineering (Lin et al., 2023). However, these approaches treat spatial reasoning as an emergent byproduct instead of an explicit modeling objective: they neither represent spatial constraints explicitly nor enforce geometric or relational consistency at the element or canvas level. Multimodal extensions using visual–language models (Ranasinghe et al., 2024; Cheng et al., 2024) improve visual grounding but still struggle with structural consistency and geometric plausibility. As illustrated in Fig. 1, GPT-5 exhibits systematic layout errors including misalignment, inconsistent sizing and violated semantic relations (e.g., underlay–text), underscoring that effective layout design requires explicit reasoning over spatial constraints and relational geometry, not merely visual grounding. Moreover, multimodal architectures often rely on inductive biases to implicitly encode geometry. A text-only formulation removes these shortcuts, making spatial reasoning an explicit and learnable capability.
These observations motivates a new research question: How can an LLM learn robust and generalizable spatial reasoning for layout design from textual serialization without pixel-level supervision? Addressing this question requires moving beyond prompt-based generation and multimodal shortcuts. However, LLMs are known to struggle with spatial reasoning, particularly in tasks involving relative geometry and multi-object relationships (Bang et al., 2023; Zha et al., 2025). In layout design, these limitations are exacerbated by (1) the lack of intrinsic geometric reasoning capability for alignment, containment and hierarchy, and (2) open-ended design spaces with sparse supervision, where multiple layouts may be equally valid and paired annotations are scarce.
To address these challenges, we propose LaySPA, a framework that equips LLMs with explicit spatial reasoning for content-aware layout design. We formulate layout generation as a policy learning problem in a spatially grounded environment, where an LLM optimizes design decisions through multi-objective spatial critique. By decoupling spatial reasoning from pixel-level generation, LaySPA enables interpretable decision making, geometry-driven generalization, and direct optimization of spatial competence via reward design
Our contributions are threefold: (1) We reformulate content-aware layout design as a policy learning problem, enabling explicit and interpretable spatial decision making beyond pixel-level generation. (2) We reinforce LLMs’ spatial reasoning over serialized canvases via multi-objective critiques that encode geometric, relational, and aesthetic constraints as learnable objectives. (3) We empirically demonstrate that LaySPA learns robust spatial reasoning, outperforming larger proprietary LLMs and most visual-based layout generators, while elucidating the complementary strengths and limitations of language-only and vision-based approaches.
2 Method
2.1 Problem Formulation
We formulate content-aware layout design as a policy learning problem, where an LLM learns a design policy $\pi_{\theta}$ that performs explicit spatial reasoning over a structured textual canvas specification $\mathcal{S}$, generating layouts $\mathcal{L}=\pi_{\theta}(\cdot\mid\mathcal{S})$ that are structurally coherent and visually appealing, without pixel-level supervision and with spatial reasoning as the primary learning signal.
Figure 2: The framework and workflow of LaySPA.
2.2 LaySPA
We propose LaySPA (Fig. 2), a reinforcement learning (§2.3) framework that grounds the canvas as a structured spatial environment (§2.2.1), produces interpretable layout actions (§2.2.2), and optimizes the LLM’s design policy via a multi-objective critique (§2.2.3).
2.2.1 Spatially-Grounded Environment
LaySPA represents each canvas as a spatially grounded environment that encodes normalized coordinates, semantic elements (e.g., text, logos, underlays), and salient regions to be avoided. This structured abstraction exposes geometric and relational constraints directly to the LLM. Formally, a layout $\mathcal{L}$ comprises a set of elements $E$ and saliency regions $S$, where each item is defined by top-left coordinates $(x_{i},y_{i})$, width and height $(w_{i},h_{i})$, and category $c_{i}$. Conditioned on element semantics and the saliency context, the LLM infers the geometric attributes $(x_{i},y_{i},w_{i},h_{i})$ for all elements. Layouts are serialized in JSON.
2.2.2 Dual-Level Design Artifacts
Given a canvas specification $\mathcal{S}=\{E,S\}$, the LLM policy $\pi_{\theta}$ generates layouts ${\ell}_{g=1}^{G}$, each comprising: (i) an explicit design reasoning trace that articulates the model’s spatial reasoning and design logic, and (ii) a structured layout specification that encodes precise geometric configurations of all elements. These outputs are specialized in $\langle design\rangle$ and $\langle layout\rangle$ blocks, respectively. This dual-level representation decouples relational reasoning from geometric instantiation, enabling interpretability, error diagnosis, and reward computation for downstream evaluation and optimization.
2.2.3 Multi-Objective Spatial Critique
To guide policy learning, we introduce a multi-objective spatial critique that evaluates candidate layouts and assigns normalized scalar rewards capturing geometric validity, structural coherence and visual aesthetics. This structured feedback enables the policy to differentiate fine-grained spatial trade-offs and converge toward globally coherent and visually balanced layouts.
1) Format Correctness Reward $R_{\text{format}}$
enforces strict compliance with the required design–layout output schema, ensuring that each output is structurally complete and parsable. Specifically, outputs must contain a $\langle\textit{design}\rangle$ block that articulates design rationale and a $\langle\textit{layout}\rangle$ block that specifies element geometries in valid JSON. The reward is assigned hierarchically: 0.1 if either block is missing, 0.2 if the JSON is unparsable, 0.5 if the JSON is valid but element types or counts mismatch the input, and 1.0 if both blocks are present, the JSON is valid, and all element types and counts match the specification. $R_{\text{format}}$ mitigates hallucinated or malformed outputs and guarantees compatibility with downstream geometric evaluation.
2) Layout Quality Reward $R_{\text{quality}}$
evaluates layouts based on geometric constraints and design principles using five normalized sub-metrics.
2.1) Inverse Collision Rate $S_{\text{icr}}$ penalizes unintended overlaps between incompatible elements (e.g., text, logos, embellishments and salient regions), while permitting intended overlaps between underlays and their associated text. Each element is represented as an axis-aligned bounding box $b_{i}=(x_{i},y_{i},w_{i},h_{i})$. For any pair $(b_{i},b_{j})$, the pairwise inverse collision score is defined as:
$S_{\text{icr}}(b_{i},b_{j})=1-\frac{\mathrm{Area}(b_{i}\cap b_{j})}{\mathrm{Area}(b_{i})+\mathrm{Area}(b_{j})-\mathrm{Area}(b_{i}\cap b_{j})}$ (1)
where $\mathrm{Area}(\cdot)$ denotes the area of a bounding box. Higher scores correspond to layouts with minimal unintended overlap.
2.2) Alignment Score $S_{\text{align}}$ quantifies layout regularity by measuring (i) global centering relative to the canvas, and (ii) mutual alignment among elements. Well-aligned layouts exhibit small deviations from the canvas center and low dispersion of element centers. Consider a layout with $N$ elements, the center of element $e_{i}$ $\boldsymbol{\mu}_{i}=\big(x_{i}+\tfrac{w_{i}}{2},\;y_{i}+\tfrac{h_{i}}{2}\big)$, and the canvas center $\boldsymbol{\mu}_{canvas}=\big(\tfrac{w_{canvas}}{2},\;\tfrac{h_{canvas}}{2}\big)$. Then we measure:
(i) Element-to-canvas alignment: valuating global centering by the normalized average Euclidean distance between element and canvas center:
$\vskip-6.99997ptA_{\text{ele-canvas}}=1-\frac{1}{N}\sum_{i=1}^{N}\frac{\|\mu_{i}-\mu_{\text{canvas}}\|_{2}}{D_{\max}}$ (2)
where $D_{\max}=\sqrt{w_{c}^{2}+h_{c}^{2}}$ ensures scale invariance.
(ii) Element-to-element alignment: capturing mutual alignment by penalizing the variance of element centers along the horizontal and vertical axes:
$A_{\text{ele-ele}}=1-\frac{1}{2}\Big(\mathrm{Var}(\mu_{i,x})+\mathrm{Var}(\mu_{i,y})\Big)\vskip-6.00006pt$ (3)
where $\text{Var}(\cdot)$ is computed across all element centers along each axis. Lower variance corresponds to higher alignment.
The final alignment score is computed as:
$S_{\text{align}}=\alpha\cdot\,A_{\text{ele-canvas}}+(1-\alpha)\cdot A_{\text{ele-ele}}\vskip-6.00006pt$ (4)
where $\alpha\in[0,1]$ balances between global centering and inter-element alignment.
2.3) Distribution Score $S_{\text{dist}}$ measures how evenly elements are arranged on the canvas by jointly evaluating spatial dispersion and area coverage. It combines (i) a normalized spread variance that encourages broad spatial distribution, and (ii) a grid coverage score that promotes utilization of the canvas while discouraging element clustering.
The normalized spread variance is defined as:
$V_{\text{spread}}=\frac{1}{N}\sum_{i=1}^{N}\frac{\lVert\boldsymbol{\mu}_{i}-\bar{\boldsymbol{\mu}}\rVert_{2}^{2}}{D_{\max}^{2}}\vskip-5.0pt$ (5)
where $\bar{\boldsymbol{\mu}}=(\bar{x},\bar{y})=\tfrac{1}{N}\sum_{i=1}^{N}\boldsymbol{\mu}_{i}$ is the average center of all elements. Higher values indicate broader spatial dispersion.
Grid coverage measures area utilization by partitioning the canvas into a $3\times 3$ grid. Let $G_{ij}\in\{0,1\}$ indicate whether grid cell $(i,j)$ contains at least one element center. The coverage score is:
$V_{\text{grid}}=\frac{1}{9}\sum_{i=1}^{3}\sum_{j=1}^{3}G_{ij}\vskip-6.00006pt$ (6)
The final distribution score is computed by:
$S_{\text{dist}}=\frac{1}{2}\big(V_{\text{spread}}+V_{\text{grid}}\big)\vskip-6.00006pt$ (7)
2.4) Spacing Consistency Score $S_{\text{spacing}}$ quantifies vertical rhythm by computing the normalized variance of distances between adjacent element centers. Elements are first sorted by vertical positions, then the distances between adjacent centers are computed. The variance of these spacings is normalized by the mean spacing and canvas height. Higher scores correspond to more even spacing. Formally, let $c_{i}^{y}=y_{i}+h_{i}/2$ denote the vertical center of element $e_{i}$. After sorting elements by $c_{i}^{y}$, the adjacent spacings are $s_{i}=c_{(i+1)}^{y}-c_{(i)}^{y}$. The spacing consistency is defined as:
$S_{\text{spacing}}=1-\frac{\text{Var}(s_{1},\dots,s_{N-1})}{\bar{s}^{2}}\vskip-3.00003pt$ (8)
where $\bar{s}$ is the average spacing. Larger values indicate more uniform vertical spacing.
2.5) Underlay-Text Constraint Score $S_{\text{underlay}}$ enforces semantic consistency by aligning each underlay with exactly one corresponding text element. Layouts are penalized if multiple text elements overlap a single underlay or if the underlay’s position and size deviate from its associated text. Exact alignment receives a score of 1, partial matches receives intermediate scores, and incorrect pairings are scored 0. This mechanism promotes readability and preserves intended functional relationships between decorative and content elements.
Finally, these sub-rewards are aggregated into a weighted layout quality score. Appendix. A Figure 6 illustrates how these scores guide the model in balancing structural feasibility and visual appeal in content-aware layouts.
3) IoU Matching Reward $R_{\text{IoU}}$
provides external supervision by measuring the geometric agreement between generated layouts and human-designed references. For each element, IoU is computed as the ratio of the intersection to the union area of the predicted and ground-truth bounding boxes. Unlike intrinsic metrics such as collision or alignment, which assess internal structural consistency, $R_{\text{IoU}}$ anchors learning to human layout exemplars that implicitly encode balance, proportion, and functional grouping. This signal guides the policy toward realistic spatial organization and visually coherent designs.
4) Final Hybrid Reward
for each generated layout is a weighted combination of:
$R(L_{i})=\lambda_{f}R_{format}+\lambda_{q}R_{quality}+\lambda_{u}R_{IoU}\vskip-5.0pt$ (9)
2.3 Training Procedure
We update the design policy using Group Relative Policy Optimization (GRPO) (Shao et al., 2024), which uses group-wise comparisons as relative baselines. At each step, the old policy $\pi_{\theta_{\text{old}}}$ samples a group of candidate layouts $G$, each scored $R(\ell_{i})$ by the hybrid reward in Eq. 9. Relative advantages are computed by normalizing rewards within the group $A(\ell_{i})=\frac{R(\ell_{i})-mean_{\ell_{i}\in G}R(\ell_{i})}{std_{\ell_{i}\in G}R(\ell_{i})}$, providing stable learning signals without an explicit critic. The policy $\pi_{\theta}$ is then optimized using the clipped surrogate objective:
$\mathcal{L}(\theta)$ $=\mathbb{E}_{\ell_{i}\sim\pi_{\theta}}\!\Big[\min\Big(\Gamma_{\theta}(\ell_{i})A(\ell_{i}),\mathrm{clip}\!\big(r_{\theta}(\ell_{i}),1-\epsilon,1+\epsilon\big)A(\ell_{i})\Big)$ (10)
$\quad-\beta\,\mathrm{KL}\!\left(\frac{\pi_{\theta}}{\pi_{\theta_{\mathrm{ref}}}}\right)\Big]$
where $\Gamma_{\theta}(\ell_{i})=\frac{\pi_{\theta}(\ell_{i})}{\pi_{\theta_{\mathrm{old}}}(\ell_{i})}$. This group-relative update prioritizes higher-reward layouts while stabilizing learning across diverse candidates.
3 Experiments
Our experiments address three questions: (Q1) Does explicit reward-driven policy optimization enable LaySPA to learn spatial reasoning beyond prompt-based LLM baselines? (Q2) How does LaySPA compare with autoregressive and multimodal layout generators in structural validity and visual quality? (Q3) What performance gap remains between text-based, policy-optimized layout design and specialized task-specific generators?
Datasets
We evaluate on CGL Zhou et al. (2022) and PKU-PosterLayout Hsu et al. (2023). PKU contains text, logo, and underlay elements, while CGL additionally includes embellishments (More details are on Appendix B). We train on 3,000 randomly sampled layouts per dataset and evaluate on full 1000 CGL and 905 PKU test instances.
Experimental Setup
LaySPA uses Qwen-2.5-7B-Instruct Team (2024) as the backbone, fine-tuned via LoRA Hu et al. (2022). Salient regions are detected using off-the-shelf models Qin et al. (2019) and converted to bounding boxes. Element geometries are masked with a [MASK] token, prompting the LLM to predict their positions $(x_{i},y_{i})$ and sizes $(w_{i},h_{i})$.
Baselines
We evaluate LaySPA against: (i) autoregressive DS-GAN (Hsu et al., 2023), (ii) VLM-based PosterLlama Seol et al. (2024), (iii) Qwen-2.5-7B-Instruct in a zero-shot setting, and (iv) proprietary GPT-4o Hurst et al. (2024). (More details in Appendix.D)
Evaluation Metrics
[…]
4 Results and Analysis
4.1 Quantitative Results
Learning Effectiveness
Figure 3 compares the base Qwen model with its LaySPA-fine-tuned counterpart on the CGL and PKU benchmarks. LaySPA delivers consistent and substantial gains across all metrics, including alignment ($63\%$), spacing consistency ($73\%$), format correctness ($14\%$) and distribution consistency ($26\%$), together with a reduction in collision rate ($36\%$). These gains indicate that reward-driven optimization enables the model to internalize explicit geometric and relational constraints, yielding layouts that better satisfy structural requirements. The largest gains obtained in alignment and spacing suggest that LaySPA particularly strengthens global spatial organization and inter-element uniformity.
((a)) CGL Dataset
((b)) PKU Dataset
Figure 3: Comparison of Qwen-7B w/wo LaySPA on CGL and PKU. Red highlights improvements.
Compared with SOTA
Table 1 reports results on CGL and PKU by different methods. PosterLlama achieves the strongest overall performance, benefiting from its visual encoders and task-specific architectural priors tailored to layout generation. LaySPA ranks second, substantially improving over its base model, with pronounced reductions in overlap and occlusion and improved underlay effectiveness, indicating more accurate functional grouping and spatial allocation. Despite larger scale, GPT-4o performs moderately, suggesting that model size alone does not guarantee fine-grained spatial reasoning in layout tasks. Notably, LaySPA attains competitive performance using only 3000 training samples per dataset, far fewer than vision-centric methods, demonstrating the strong spatial inductive bias and sample efficiency induced by reward-driven policy optimization.
Table 1: Comparison of overlap (Ove↓), underlay effectiveness (Und↑) and occlusion (Occ↓) on CGL and PKU datasets. “Real Data" denotes the annotated layout. Best results are bolded; second-best are underlined.
#Params CGL Dataset PKU Dataset
Model Ove ↓ Und ↑ Occ ↓ Ove ↓ Und ↑ Occ ↓
Real Data - 0.0003 0.9926 0.1379 0.0013 0.9974 0.1828
DS-GAN 30M 0.0361 0.6309 0.1521 0.0336 0.7613 0.2574
PosterLLama 7B 0.0024 0.9918 0.1476 0.0032 0.9998 0.2087
GPT-4o 200B 0.0365 0.5873 0.1591 0.0371 0.6384 0.2743
Qwen-7B 7B 0.0474 0.5729 0.1615 0.0479 0.6059 0.2384
Qwen-7B++LaySPA 7B 0.0257 0.6989 0.1487 0.0260 0.7688 0.2072
4.2 Qualitative Results
Figure 4 presents qualitative comparisons across methods. DS-GAN achieves reasonable canvas coverage and consistent element sizing but often exhibits misalignment, unintended overlaps, and interference with salient regions, reflecting limited global spatial coordination. PosterLlama produces the most visually coherent layouts, benefiting from vision-specific encoders that impose strong structural priors and balanced compositions. GPT-4o underperforms particularly on CGL, with frequent misalignment and weak relational organization, revealing limitations in fine-grained spatial control despite its scale. In contrast, LaySPA consistently generates visually harmonious and well-proportioned layouts, demonstrating that explicitly learned spatial reasoning and policy optimization improves overall design quality.
((a))
((b))
Figure 4: Layouts generated by different methods.
4.3 Ablation Study
To dissect the contributions of each component in the hybrid reward (Eq. 9), we vary $(\lambda_{f},\lambda_{q},\lambda_{u})$ and evaluate LaySPA on the CGL benchmark. We consider four configurations: (1) Format-focused $(0.5,0.4,0.1)$ emphasizes schema compliance, (2) Quality-focused $(0.1,0.8,0.1)$ prioritizes geometric and visual design principles, (3) IoU-focused $(0.1,0.1,0.8)$ emphasizes alignment with reference layouts, and (4) Balanced-hybrid $(0.1,0.45,0.45)$ distributes weight across all objectives.
Figure 5 shows performance across collision, alignment, spacing consistency, and distribution metrics. Quality-focused training yields the best overall results, highlighting the importance of explicit geometric and relational objectives for coherent, functionally valid layouts. The balanced-hybrid setting offers a pragmatic trade-off, maintaining reasonable reference alignment while preserving structural fidelity. In contrast, IoU-focused underperforms, indicating that optimizing for reference matching alone leads to brittle spatial reasoning and insufficient attention to geometric constraints.
Figure 5: Ablation study on varying reward weights
5 Related Works
Spatial Reasoning with LLMs
Vision–Language Models (VLMs) (Fan et al., 2025) and Multimodal LLMs (MLLMs) (Wu et al., 2025a) enable spatial reasoning by grounding language in visual inputs, using multi-view images (Qi12 et al., ; Fu et al., 2025; Gu et al., 2024), depth cues (Ma et al., 2024; Cheng et al., 2024; Cai et al., 2025) or cross-modal embeddings (Lüddecke and Ecker, 2022; Chen et al., 2024). While effective, these methods are computationally intensive, encode geometry implicitly, and offer limited interpretability.
Content-Aware Layout Generation
Traditional layout generation relies on deep generative models such as VAEs (Jyothi et al., 2019), GANs (Zhou et al., 2022; Hsu et al., 2023; Li et al., 2019), transformers (Kong et al., 2022), autoregressive (Horita et al., 2024), and diffusion-based models (Chai et al., 2023a, b), supplemented by retrieval (Horita et al., 2024) or post-hoc refiners (Shen et al., 2025) to reduce overlap and misalignment. These approaches rely heavily on annotated data and thus generalizing poorly beyond training distributions. LaySPA differs by formulating layout design as reinforcement learning over a structured textual environment, enabling diverse exploration of geometrically valid layouts guided by spatial objectives.
LLM-based Layout Generation
Recent works leverage LLMs’ design priors by serializing layouts into text formats such as HTML (Tang et al., 2023; Seol et al., 2024), CSS (Feng et al., 2023), or SVG (Hsu and Peng, 2025) and applying recursive reasoning to capture spatial dependencies (Tian et al., 2025). Retrieval-based pipelines (Lin et al., 2023; Forouzandehmehr et al., 2025) ground outputs in exemplar layouts and iteratively refine them using LLMs and vision–language feedback. These methods often rely on emergent capabilities and exemplar retrieval, whereas LaySPA is explicitly guided by spatial and visual rewards and optimized through peer-relative policy learning.
6 Conclusion
We introduce LaySPA, a GRPO-based framework that reframes content-aware layout design as a policy learning problem, guiding LLMs to perform explicit spatial reasoning over structured textual environments. Optimized with multi-objective spatial critiques, LaySPA learns robust and generalizable spatial behaviors for layout design. Empirically, LaySPA substantially improves layout quality, outperforms larger proprietary LLMs, and approaches specialized multimodal generators without visual inputs and less training samples. These results demonstrate that spatial reasoning can be explicitly optimized in text, positioning language-grounded, reward-driven policy learning as a scalable alternative to vision-centric layout design. Limitations include reliance on precomputed saliency and single-step generation. Future work will explore richer semantic grounding, multi-turn reinforcement learning, and hierarchical design policies.
Acknowledgments
Original images from PKU-PosterLayout dataset by Hsu et al. available at https://huggingface.co/datasets/creative-graphic-design/PKU-PosterLayout, licensed under CC BY-SA 4.0. Original images from CGL dataset by Zhou et al. available at https://huggingface.co/datasets/creative-graphic-design/CGL-Dataset, licensed under CC BY-NC-SA 4.0.
Appendix A Visualization of Rewards
Figure 6 illustrates how individual components of the layout quality score guide the agent toward content-aware designs that jointly satisfy structural feasibility and visual appeal.
Collision $R_{{icr}}$ Refer to caption Refer to caption
Alignment $R_{{al}}$ Refer to caption Refer to caption
Distribution $R_{{dis}}$ Refer to caption Refer to caption
Spacing $R_{{sp}}$ Refer to caption Refer to caption
Underlay-text $R_{{ut}}$ Refer to caption Refer to caption
Figure 6: An illustration of how each layout quality score functions guides the agent toward human-preferred designs () while discouraging () undesirable spatial arrangements.
Appendix B Dataset
Table 2 lists the details of the CGL and PKU benchmark.
Table 2: Layout generation datasets.
Dataset CGL PKU
# Layouts 60,548 9,974
# Canvases 1,000 905
Elements Text, Logo, Underlay, Emb. Text, Logo, Underlay
Canvas Non-empty Non-empty
Categories Cosmetics, Elec., Clothing, etc. Cosmetics, Elec., Clothing, Food, Toys, etc.
Appendix C Prompts
The followings are the prompts used in the fine-tuning and inference.
User Prompt <task> You are a content-aware graphic layout designer. Given a canvas with fixed width and height, along with layout elements and saliency regions specified in JSON format, your task is to perform explicit spatial reasoning over the structured representation and predict each element’s geometry, top-left coordinates and size $(x,y,w,h)$ to produce a coherent and visually balanced layout. </task> <objective> Design a coherent, balanced, and visually appealing layout that satisfies all structural and spatial constraints. </objective> <instructions> **Saliency awareness.** Saliency boxes denote visually important regions; no element may overlap them. **Element interactions.** Text and logo elements must not overlap. An underlay may overlap only with its associated text element. **Underlay-text coupling.** Each underlay supports at most one text element and should approximately align in position and size with that text. **Global composition.** Favor centered layouts, even spacing, clear grouping, and structural symmetry to achieve a harmonious overall design. </instructions>
User Prompt The layout to be generated should have $width={example[^{\prime}width^{\prime}]}$ and $height={example[^{\prime}height^{\prime}]}.$
Below is a JSON layout where the element locations and sizes $(top_{l}eft_{x},top_{l}eft_{y},width,height)$ are masked as [MASK].Allocate elements to positions that meet the above structural requirements and replace [MASK] with integers.
The layout is: {canvas_json}
And the saliency boxes are: {saliency_box}
The output format of reasoning and final layout are enclosed with <design></design> and <layout></layout> tags:
<design>Concise step-by-step reasoning in less than 50 words</design>
<layout>Optimized layout in the same format as the input layout</layout> )
Appendix D Baseline Methods
  • DS-GAN (Hsu et al., 2023): DS-GAN is a CNN-LSTM-based conditional generative adversarial network that learns design sequences conditioned on a given canvas to automatically generate content-aware visual-textual presentation layouts.
  • PosterLlama (Seol et al., 2024): a multimodal layout generation system that uses CodeLLaMA-7B as the language backbone with a DINOv2 visual encoder. Layouts are reformulated as HTML sequences. The model is trained via supervised instruction tuning with depth-based data augmentation.
  • Qwen-2.5-7B-Instruct: A dense transformer language model with 7 billion parameters. Given the same inputs as LaySPA, while used in a training-free, prompt-based setting to generate geometric attributes for all layout elements.
  • GPT-4o: A large-scale proprietary multimodal foundation model that takes the canvas as visual input and is prompted to generate the geometric specifications of all layout elements.